### v1

@@ V-A Experimental Results
Figure 13: Performance metrics for the virtual welding task across four TOIs for (a) quiet stance and (b) kneeling gaits.
Fig. 12 further compares the sway acceleration ellipses and the postural sway ellipses for TOIs 2 and 4 experiments. The relatively larger ellipse areas for the kneeling gait compared to the stance gait, particularly under TOI 2, suggest that kneeling is inherently less stable than quiet stance. In the presence of higher cognitive load and visual disturbances, the exoskeleton provides significant support and stabilization during kneeling tasks, as evidenced by the reduced ellipse size compared to TOI 2. Table III lists the mean and standard deviation of $A_{t}^{\text{Acc}}$ and $A_{t}^{\text{Sway}}$ as quantitative assessments of postural stability. For quiet stance, there is a notable decrease in average $A_{t}^{\text{Sway}}$ from TOI 1 and 2, signaling that subjects reduced their sway in response to visually induced stress. However, an increase of $A_{t}^{\text{Acc}}$ suggests that the subjects responded with significant dynamic control to maintain this constrained posture. The stabilizing effect of the exoskeleton was evident in TOI 3, where reduced $A_{t}^{\text{Sway}}$ and $A_{t}^{\text{Acc}}$ correspond to a relaxed stance, while in TOI 4, with higher $A_{t}^{\text{Sway}}$ but lower $A_{t}^{\text{Acc}}$ values, indicates deliberate and controlled movements. Welding tasks, particularly TOI 6, show the largest $A_{t}^{\text{Acc}}$ and $A_{t}^{\text{Sway}}$, indicating that engagement in the task necessitates broader movements. For kneeling condition, the overall $A_{d}^{\text{Acc}}$ and $A_{d}^{\text{Sway}}$ are smaller compared to the quiet stance condition, suggesting that the kneeling posture inherently constrains motion.
@@
Fig. 13 shows the assessment of task performance through three metrics for the welding task across TOI 5 to 8. For quiet stance, it is observed that the incorporation of the exoskeleton in TOIs 7 and 8 correlates with improved accuracy and completion rate. The completion rate in HE conditions (i.e., TOI 8) is particularly enhanced with exoskeleton assistance, whereas precision appears to be less influenced by the elevation or the presence of the exoskeleton. The kneeling gait exhibits a similar trend in task performance metrics. The use of the exoskeleton in TOIs 7 and 8 under kneeling conditions also leads to increased accuracy and completion rate compared to TOIs 5 and 6 without exoskeleton assistance. This finding suggests that the exoskeleton’s benefits extend beyond a quiet stance and can effectively support task execution in kneeling postures as well. However, it is worth noting that the overall accuracy and completion rate values for kneeling are slightly lower than those for quiet stance across all TOIs. This difference may be attributed to the inherent challenges associated with maintaining balance and performing tasks in a kneeling posture.
TABLE V: Subject Questionnaire Results
Subject ID S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11
CA Quest. Fear Heights $22$ $56$ $40$ $40$ $27$ $53$ $54$ $33$ $48$ $45$ $60$
Avoid. HE $22$ $36$ $26$ $28$ $23$ $36$ $34$ $27$ $49$ $36$ $37$
Cog. Load Stressful HE+Exo HE+Exo HE HE HE HE/HE+Exo HE HE HE HE HE
Least Focus LE+Exo LE HE HE HE LE+Exo/LE LE HE HE HE HE
@@
Fig. 14 shows the heart rate data collected during the welding tasks for both gaits. During quiet stance, the heart rate increased from LE to HE suggests a progressive elevation in stress levels, with the highest median heart rate recorded during the HE tasks, regardless of exoskeleton usage. This incremental pattern reflects the physiological demands imposed by elevation changes and the complexity of tasks performed under such conditions. With the introduction of the knee exoskeleton, subjects experienced a reduction in stress levels during tasks at both LE and HE. During kneeling gaits, although the application of the exoskeleton reduced human effort and lowered the heart rate, the heart rate was generally higher compared to a quiet stance. Table IV lists the average heart rates from some subjects for TOI 5 to 8. The data reveal minimal fluctuations in heart rate among most subjects across different TOIs, suggesting that the welding task may serve as a focal stimulus, concentrating the participants’ attention and thereby stabilizing physiological responses [23]. Subjects S2, S3, and S5 demonstrated notable variations in average heart rate upon the integration of the exoskeleton into the task during stance gait. Similar trends are observed during kneeling gaits, where most subjects experienced a reduction in heart rates with exoskeleton assistance while performing welding tasks.
@@
Table V summarizes the results from the subjects’ questionnaire feedback. The data indicates a wide range of susceptibility to fear and acrophobia across all subjects for both quiet stance and kneeling gaits. In quiet stance experiments, subjects S2 and S7 showed the highest levels of susceptibility to fear and acrophobia, whereas S1 displayed the lowest. For kneeling tests, S11 exhibited the highest susceptibility to fear and S9 had the lowest. Cognitive loading was reported to be higher during tasks at HE for most subjects than that at LE in both stance and kneeling experiments. However, subjects S1 and S2 in quiet stance and S6 in kneeling experiments reported high cognitive load with the exoskeletons at HE. When considering task performance, subjects generally focused best on the addition of the exoskeletons, while their focus was compromised without the exoskeletons. This trend was observed in both gaits experiments, suggesting that the exoskeleton may offer a psychological benefit in high-stress situations, regardless of the working posture.
All subjects reported feeling tense during HE scenarios and calm during low elevation scenarios, with varied responses to the presence of the exoskeleton. However, it should be noted that the subjects’ perception of focus and relaxation does not always equate to good postural balance and balancing choices. Stressful situations, such as focusing on welding, can lead to instinctive balance strategies being chosen, resulting in better balance. Perturbations under stressful situations can cause posture to rapidly become unstable. In the stance experiments, subjects S1 and S2 expressed an increase in stability and balance when utilizing the exoskeleton. In contrast, S7 experienced a restriction in mobility due to the exoskeleton. Meanwhile, S5 noticed a marginal benefit, and S6 found that the exoskeletons helped maintain position during the task. The responses to height exposure were equally varied. Most subjects consistently described the HE environment as highly realistic and fear-inducing. Interestingly, all participants in the kneeling experiments reported enhanced stability and reduced effort in maintaining balance when using the exoskeletons.
V-B Simulation Results
The best-fit parameters were derived by varying the parameter sets to obtain frequency behavior matching the observed frequency behavior of humans during quiet stance across various TOIs. Fig. 15 demonstrates an example of exponential fitting on experimental stance and kneeling IP curves. Table VI lists the resulting best-fit parameter sets that quantify the favored balance strategy corresponding to each TOI. The simulated IP curve that generates the minimum error in comparison to the exponentially fitted experimental IP curve is identified.
@@ V-B Simulation Results
Examining the variations in ${\alpha}$ values reveals a characteristic for minimal control effort. A high ${\alpha}$ magnitude signifies a minimal control effort and is generally chosen by humans during quiet stance, yet this tendency differs across different TOIs. For the balancing experiments, all cases had a high magnitude of ${\alpha}$, with TOI 2 utilizing the lowest control effort. The same results were observed from the task performance scenarios TOIs 5, 6, and 8, while TOI 7 had a lower magnitude of ${\alpha}=10$. The low value of ${\alpha}$ implies that a non-minimal effort control strategy was unexpectedly chosen to balance. This anomaly potentially suggests that the subjects could have altered their balance strategies to adapt to the exoskeleton in the presence of non-threatening visuals, rather than passively rely on the extra support it provides. For the rest of the welding cases, ${\alpha}$ value (i.e., $10^{10}$) is high, which implies that the welder’s main focus on precision during the welding task caused them to use a minimal effort balance strategy.
The balance strategies were further refined by the choice of joint strategy, as determined by the ${\boldsymbol{\beta}}$ values. Across all TOIs, it was found that $\beta_{2}<\beta_{1}<\beta_{3}$, which suggests employing hip strategy is penalized compared to employing ankle-knee strategies. TOIs 2 and 7 have higher $\beta_{3}$ values, with subjects favoring the ankle-knee strategy heavily over the hip strategy to balance. The preference for the ankle-knee strategy in TOI 2 can be the consequence of the destabilizing visuals, with the subjects attempting to keep their head, the visual source of the body, stable and upright. For potentially similar reasons, TOI 6 shows relatively higher noise values at the ankle ($\sigma_{1}$) and knee joints ($\sigma_{2}$) than at the hip joint ($\sigma_{3}$). For TOI 7, the high effort control strategy suggests that the ankle-knee strategy chosen, while intentional, required much effort to maintain balance. The neuromechanical noise at the knee joint is lower with the presence of the exoskeleton for the stance case (TOIs 3 and 4). On introducing the VR task scenario, the neuromechanical noise experienced is similar at all joints with the exoskeleton. A similar balance strategy is used for both TOI 3 and 4, thus exhibiting reduced influence of visual stimuli on balance with the exoskeleton in the absence of task performance.
@@ VI Discussion
The balance strategies discerned from IP curve analyses suggest active neuromuscular engagement at lower frequencies and a transition to passive control strategies at higher frequencies, modulated by the knee exoskeleton. The decrease in mean CF $\bar{\omega}^{\text{CF}}$ with exoskeleton application implies an adaptive response in the human balance control system, possibly reflecting augmented stability. The consistency in the mean HFA slope $\bar{k}^{\text{HFA}}$ across TOIs with exoskeleton intervention indicates uniform neuromechanical impedance on the joints, with the high slope interpreted as an increment in neuromechanical noise and a strategic shift in balance at high frequencies. The altered strategy can be a consequence of the individuals’ effort to recalibrate their balance strategy to accommodate the exoskeleton during task execution (e.g., TOI 7), as opposed to being passively supported in a static posture (TOI 8). The results from Tables II align with the notion in [27] that while exoskeletons can enhance postural control, their impact on the neuromechanical behavior of the body is predictable and can be characterized by specific frequency-domain features of the IP curve. These insights are pivotal for optimizing the design of exoskeletons to harmonize with the body’s natural balance strategies. By observing the parameter $\alpha$, it was also found that kneeling balance control requires much higher effort than during stance to maintain postural stability.
@@
As shown in Fig. 7, the introduction of the welding task increased presence of low frequency spikes in the rest of the subjects’ IP curves. Low frequency spikes reflect proactive intentional efforts to change the balance strategy rather than reactionary attempts to stabilize stance in response to environmental perturbations. While low frequency spikes reduced with exoskeleton assistance for S5, they increased for S6 for both balance and task performance TOIs. The subject specific insights drawn from the IP analysis are valuable in identifying subjects who may be susceptible to destabilization due to specific task scenarios, thereby informing necessary measures to mitigate the risk of injury.
@@
Task performance assessments reveal the exoskeleton’s potential to alleviate the cognitive and physical loads on the subject in the presence of higher cognitive load and visual disturbances in construction. The comparative analysis between quiet stance and kneeling emphasizes the need for targeted interventions and training programs addressing specific challenges associated with different working postures. The increase in heart rate during kneeling can be attributed to greater muscle activation, reduced venous return, and psychological stress [36, 37]. The combination of physiological metrics permits a distinction between physical and cognitive stressors, highlighting the complexity of the exoskeleton’s effectiveness in high-stress tasks. The consistency in precision values suggest that the exoskeleton primarily aids in postural stability and task completion rather than fine motor control.
@@
By considering the GRF intersection point, the IP method provides a comprehensive assessment of the worker’s postural control mechanisms. This approach is particularly valuable in the context of construction work, where workers are exposed to a wide range of environmental conditions, such as elevated work surfaces and unstable footing. By analyzing the frequency characteristics of the IP heights, researchers and safety professionals can gain insights into how workers adapt their balance strategies in response to different work scenarios. It is noteworthy that only a specific subset of ${\boldsymbol{\beta}}$ value sets could effectively describe the IP characteristics across various TOIs. The knee joint strategies play a crucial and non-trivial role in maintaining balance, highlighting the importance of considering knee joint control when designing balance-assistive devices or developing balance training programs for construction workers.
In contrast to the previous work in [6], which focused on a single-link inverted pendulum model, the current study employs a more comprehensive approach to capture the complexity of balance control in real-world construction scenarios. The research work in [6] employed the PSD analysis to examine postural balance. This metric, as demonstrated in this study, can providing valuable insights into the dominant frequencies of sway. However, PSD analysis alone cannot provide a comprehensive understanding of the underlying neuromechanical strategies employed by individuals to maintain balance (see Fig. 9), as it does not capture the spatial organization, coordination, or phase relationships between different body segments. Using the IP, multi-link inverted pendulum models and the LQR controller, the current work analyzed the model’s response to various conditions and the interpretation of the results in terms of physiological and biomechanical implications. The LQR controller aligns with the hypothesis that the human neural system employs an optimal control strategy to maintain balance [25, 26] and has been supported by numerous studies (e.g., [38]).
@@
Despite the reported benefits in terms of stability and balance, some subjects expressed that wearing the knee exoskeleton was less comfortable during kneeling compared to the stance posture. This disparity in comfort levels may be attributed to the unique biomechanical demands and joint configurations associated with kneeling. These observations underscore the importance of considering posture-specific design factors and adaptability when developing wearable exoskeletons for construction workers. There are some other limitations to consider for future research directions. First, although the VR environment allows for a controlled and safe investigation of various scenarios, future work should further validate these findings in real-world construction settings. Second, we recruited young and healthy subjects for this study and we should expand the subject pool to include experienced construction workers of all ages to provide a comprehensive evaluation of the exoskeletons’ impact on balance and performance.

### v2

@@ V Results
Fig. 12 further compares the sway acceleration ellipses and the postural sway ellipses for TOIs 2 and 4 experiments. The relatively larger ellipse areas for the kneeling gait compared to the stance gait, particularly under TOI 2, suggest that kneeling is inherently less stable than stance. In the presence of higher cognitive load and visual disturbances, the exoskeleton provides significant support and stabilization during kneeling tasks, as evidenced by the reduced ellipse size compared to TOI 2. Table III lists the mean and standard deviation of $A_{t}^{\text{Acc}}$ and $A_{t}^{\text{Sway}}$ as quantitative assessments of postural stability. For stance, there is a notable decrease in average $A_{t}^{\text{Sway}}$ from TOI 1 and 2, signaling that subjects reduced their sway in response to visually induced stress. However, an increase of $A_{t}^{\text{Acc}}$ suggests that the subjects responded with significant dynamic control to maintain this constrained posture. The stabilizing effect of the exoskeleton was evident in TOI 3, where reduced $A_{t}^{\text{Sway}}$ and $A_{t}^{\text{Acc}}$ correspond to a relaxed stance, while in TOI 4, with high $A_{t}^{\text{Sway}}$ but low $A_{t}^{\text{Acc}}$ values, indicates deliberate and controlled movements. Welding tasks, particularly TOI 6, show the largest $A_{t}^{\text{Acc}}$ and $A_{t}^{\text{Sway}}$, indicating that engagement in the task necessitates broader movements. For the kneeling condition, the overall $A_{d}^{\text{Acc}}$ and $A_{d}^{\text{Sway}}$ are smaller compared to the quiet stance condition, suggesting that the kneeling posture inherently constrains motion.
@@
Figure 13: Performance metrics for the virtual welding task across four TOIs for (a) quiet stance and (b) kneeling gaits. Error bars represent ±$\pm 1$ standard error. Asterisks indicate significant differences between the two groups ($p<0.05$).
@@
Fig. 13 shows the task performance through three metrics for the welding task across TOI 5 to 8. The ANOVA analysis revealed significant differences across TOIs for performance metrics. For quiet stance, it is observed that the incorporation of the exoskeleton in TOIs 7 and 8 correlates with improved accuracy and completion rate. Specifically, accuracy and completion rate were significantly higher in TOI 8 compared to TOI 5 ($p<0.05$) and this suggests that the exoskeletons improve performance at high elevations. The completion rate in HE conditions (i.e., TOI 8) is particularly enhanced with exoskeleton assistance, whereas precision appears to be less influenced by the elevation or the presence of the exoskeleton. The kneeling gait exhibits a similar trend in task performance metrics. Precision was significantly higher in TOI 8 compared to TOI 5 ($p<0.05$). The use of the exoskeleton in TOIs 7 and 8 in kneeling gaits also leads to increased accuracy and completion rate compared to TOIs 5 and 6 without exoskeleton assistance. This finding suggests that the exoskeleton’s benefits extend beyond a quiet stance and effectively support task execution in kneeling postures as well. However, it is worth noting that the overall accuracy and completion rate values for kneeling are slightly lower than those for quiet stance across all TOIs. This difference may be attributed to the inherent challenges associated with maintaining balance and performing tasks in kneeling posture.
@@
Fig. 14 shows the heart rate data collected during the welding tasks for both gaits. During quiet stance, the heart rate increase from LE to HE suggests a progressive elevation in stress levels, with the highest median heart rate recorded during the HE tasks, regardless of exoskeleton usage. This incremental pattern reflects the physiological demands imposed by elevation changes and the complexity of tasks performed under such conditions. Statistical analysis using ANOVA revealed that the differences in heart rates between the conditions were not statistically significant ($p>0.05$), indicating that while heart rates increased from LE to HE, the variability among subjects may have masked any significant group-level effects. Nevertheless, the trends observed, particularly the reduction in heart rate with exoskeleton assistance, align with our hypothesis that the exoskeleton alleviates physical stress during tasks at both LE and HE. During kneeling gaits, although the application of the exoskeleton reduced human effort and lowered the heart rate, the heart rate was generally higher compared to a quiet stance. The observed trend of lower heart rates with exoskeleton usage suggests potential physiological benefits in reducing stress, even if the differences did not reach statistical significance. Table IV lists the average heart rates from some subjects for TOI 5 to 8. The data reveal minimal fluctuations in heart rate among most subjects across different TOIs, suggesting that the welding task may serve as a focal stimulus, concentrating the participants’ attention and thereby stabilizing physiological responses [23]. Subjects S2, S3, and S5 demonstrated notable variations in average heart rate upon the integration of the exoskeleton into the task during stance gait. Similar trends are observed during kneeling gaits, where most subjects experienced a reduction in heart rates with exoskeleton assistance while performing welding tasks.
@@
Table V summarizes the results from the subjects’ questionnaire feedback. The data indicates a wide range of susceptibility to fear and acrophobia across all subjects for both stance and kneeling gaits. In stance experiments, subjects S2 and S7 showed the highest levels of susceptibility to fear and acrophobia, whereas S1 displayed the lowest. For kneeling tests, S11 exhibited the highest susceptibility to fear and S9 had the lowest. Cognitive loading was reported to be higher during tasks at HE for most subjects than that at LE in both stance and kneeling experiments. However, subjects S1 and S2 in stance and S6 in kneeling experiments reported high cognitive load with the exoskeletons at HE. When considering task performance, subjects generally focused best on the addition of the exoskeletons, while their focus was compromised without the exoskeletons. This trend was observed in both gaits experiments, suggesting that the exoskeleton might offer a psychological benefit in high-stress situations, regardless of the working posture.
TABLE V: Subject Questionnaire Results
Subject ID S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11
CA Quest. Fear Heights $22$ $56$ $40$ $40$ $27$ $53$ $54$ $33$ $48$ $45$ $60$
Avoid. HE $22$ $36$ $26$ $28$ $23$ $36$ $34$ $27$ $49$ $36$ $37$
Cog. Load Stressful HE+Exo HE+Exo HE HE HE HE/HE+Exo HE HE HE HE HE
Least Focus LE+Exo LE HE HE HE LE+Exo/LE LE HE HE HE HE
All subjects reported feeling tense during HE scenarios and calm during low elevation scenarios, with varied responses to the presence of the exoskeleton. However, it should be noted that the subjects’ perception of focus and relaxation did not always equate to good postural balance and balancing choices. Stressful situations, such as focusing on welding, can lead to instinctive balance strategies being chosen, resulting in better balance. Perturbations under stressful situations can cause posture to rapidly become unstable. In the stance experiments, subjects S1 and S2 expressed an increase in stability and balance when utilizing the exoskeletons. In contrast, S7 experienced a restriction in mobility due to the exoskeletons. Meanwhile, S5 noticed a marginal benefit, and S6 found that the exoskeletons helped maintain position during the task. The responses to height exposure were equally varied. Most subjects consistently described the HE environment as highly realistic and fear-inducing. Interestingly, all participants in the kneeling experiments reported enhanced stability and reduced effort in maintaining balance when using the exoskeletons.
The best-fit parameters were derived by varying the parameter sets to obtain frequency behavior matching the observed experiments during quiet stance across various TOIs. Fig. 15 shows an example of exponential fitting the stance and kneeling IP curves. Table VI lists the best-fit parameter sets that quantify the favored balance strategy corresponding to each TOI. The parameter values were obtained by minimizing the sum of the angle $q(t)$ errors from all subject experiments.
@@
Examining the variations in ${\alpha}$ values reveals a characteristic for minimal control effort. A high ${\alpha}$ magnitude signifies a minimal control effort and is generally chosen by humans during stance and this tendency however differs across TOIs. For the balancing experiments, all cases had a high magnitude of ${\alpha}$, with TOI 2 utilizing the lowest control effort. The same results were observed from the task performance scenarios TOIs 5, 6, and 8, while TOI 7 had a low magnitude of ${\alpha}=10$. The low value of ${\alpha}$ implies that a non-minimal effort control strategy was unexpectedly chosen to balance. This anomaly potentially suggests that the subjects could have altered their balance strategies to adapt to the exoskeleton in the presence of non-threatening visuals, rather than passively rely on the extra support it provides. For the rest of the welding cases, ${\alpha}$ value (i.e., $10^{10}$) is high, which implies that the welder’s main focus on precision during the welding task caused them to use a minimal effort balance strategy.
@@
The LQR parameter ${\boldsymbol{\beta}}$ provides insights into joint-specific balance strategies and reveals how control efforts are distributed across different joints. Large $\beta_{i}$ value ($i=1,2,3$) implies a minimal control effort to the corresponding joint. Across all TOIs, it was found that $\beta_{2}<\beta_{1}<\beta_{3}$, which suggests employing hip strategy is penalized compared to employing ankle-knee strategies. TOIs 2 and 7 have higher $\beta_{3}$ values, with subjects favoring the ankle-knee strategy heavily over the hip strategy to balance. The preference for the ankle-knee strategy in TOI 2 can be the consequence of the destabilizing visuals, with the subjects attempting to keep their head, the visual source of the body, stable and upright. For possible similar reasons, TOI 6 shows relatively higher noise values at the ankle ($\sigma_{1}$) and knee joints ($\sigma_{2}$) than at the hip joint ($\sigma_{3}$). For TOI 7, the high-effort control strategy suggests that the ankle-knee strategy chosen, while intentional, required much effort to maintain balance. The neuromechanical noise at the knee joint is lower with the presence of the exoskeleton for the stance case (TOIs 3 and 4). On introducing the VR task scenario, the neuromechanical noise experienced is similar at all joints with the exoskeleton. A similar balance strategy is used for both TOI 3 and 4, thus exhibiting reduced the influence of visual stimuli on balance with the exoskeleton in the absence of task performance.
@@ VI Discussion
The balance strategies discerned from IP curve analyses suggest active neuromuscular engagement at lower frequencies and a transition to passive control strategies at higher frequencies, modulated by the used knee exoskeleton. The decrease in mean CF with exoskeleton application implies an adaptive response in the human balance control system, possibly reflecting augmented stability. The consistency in the mean HFA slope across TOIs with exoskeleton intervention indicates uniform neuromechanical impedance on the joints, with the high slope interpreted as an increment in neuromechanical noise and a strategic shift in balance at high frequencies. The altered strategy can be a consequence of the individuals’ effort to recalibrate their balance strategy to accommodate the exoskeleton during task execution (e.g., TOI 7), as opposed to being passively supported in a static posture (TOI 8). The results from Tables II align with the notion in [27] that while exoskeletons can enhance postural control, their impact on the neuromechanical behavior of the body is predictable and can be characterized by specific frequency-domain features of the IP curve. These insights are pivotal for optimizing the design of exoskeletons to harmonize with the body’s natural balance strategies. By observing the parameter $\alpha$, it is also found that kneeling balance control requires much higher effort than during stance to maintain postural stability.
@@
As shown in Fig. 7, the introduction of the welding task increased the presence of low-frequency spikes in the rest of the subjects’ IP curves. Low-frequency spikes reflect proactive intentional efforts to change the balance strategy rather than reactionary attempts to stabilize stance in response to environmental perturbations. While low-frequency spikes were reduced with exoskeleton assistance for S5, they increased for S6 for both balance and task performance TOIs. The colorbluesubject-specific insights drawn from the IP analysis are valuable in identifying subjects who may be susceptible to destabilization due to specific task scenarios, thereby informing necessary measures to mitigate the risk of injury.
@@
Task performance assessments reveal the exoskeleton’s potential to alleviate the cognitive and physical loads on the subject in the presence of visual disturbances in construction. The comparative analysis between quiet stance and kneeling emphasizes the need for targeted interventions and training programs addressing specific challenges associated with different working postures. The increase in heart rate during kneeling can be attributed to greater muscle activation, reduced venous return, and psychological stress [37, 38]. The combination of physiological metrics permits a distinction between physical and cognitive stressors, highlighting the complexity of the exoskeleton’s effectiveness in high-stress tasks. The consistency in precision values suggests that the exoskeleton primarily aid in postural stability and task completion rather than fine motor control.
@@
The IP method provides a comprehensive assessment of the worker’s postural control mechanisms. This approach is particularly valuable in the context of construction work, where workers are exposed to a wide range of environmental conditions, such as elevated worksite and unstable floor surfaces. By analyzing the frequency characteristics of the IP heights, researchers and safety professionals can gain insights into how workers adapt their balance strategies in response to different work scenarios. The knee joint strategies play a crucial and non-trivial role in maintaining balance, highlighting the importance of considering knee joint control when designing balance-assistive devices or developing balance training programs for construction workers. In contrast to the previous work in [6], the current study employs a more comprehensive approach to capture the complexity of balance control in real-world construction scenarios. The research work in [6] employed the PSD analysis to examine postural balance. This metric, as demonstrated in this study, can provide valuable insights into the dominant frequencies of sway. However, PSD analysis alone cannot provide a comprehensive understanding of the underlying neuromechanical strategies employed by individuals to maintain balance (see Fig. 9), as it does not capture the spatial organization, coordination, or phase relationships between different body segments.
Using the IP, multi-link inverted pendulum models, and the LQR controller, the current work analyzed the model’s response to various conditions and provided the interpretation in terms of physiological and biomechanical implications. The LQR controller aligns with the hypothesis of optimal control in human neural balance systems [25, 26, 39] and provides a framework to interpret the physiological mechanisms underlying postural control. These mechanisms, governed by intricate neural feedback loops and coordinated muscle activation patterns across the ankle, knee, and hip joints, are reflected in the LQR parameters. For example, small $\beta_{1}$ values during stance indicated the ankle’s dominant role in balance, suggesting heightened activation of ankle muscles and associated proprioceptive feedback systems. Conversely, large $\alpha$ values signified minimal neural and muscular activation, implying reliance on passive biomechanical properties for stability. The increase in $|\bar{\omega}^{\text{CF}}|$ under conditions such as elevated heights or visual perturbations revealed heightened neural and muscular engagement. This physiological response manifested as increased muscle co-contraction and enhanced sensory integration to maintain balance. Notably, the exoskeleton’s influence on these physiological responses was evident in the reduced $|\bar{\omega}^{\text{CF}}|$ and higher $\alpha$ values, indicating a shift towards passive control strategies. This shift suggests decreased muscle activation, particularly at the knee joint. Physiologically, this translates to reduced muscular fatigue and improved postural stability over prolonged use.
The real-world application of the findings in this study is twofold. First, the results in this work demonstrate that exoskeletons enhance balance and reduce postural sway, particularly in environments such as high elevations or unstable floor surfaces. This suggests that exoskeletons play a significant role in reducing fall risks and improving worker safety. Moreover, enhanced postural stability in skilled tasks such as welding implies that exoskeletons can also contribute to reduced work errors and lead to improved productivity. Additionally, the research outcomes provide exoskeleton designers insights into joint-level balance strategies. This helps develop personalized control algorithms that respond to a worker’s changing balance strategy and tailor the device design to individual users. The method can also be integrated into VR/MR-based training programs, allowing workers to safely practice and improve their balance strategies in a controlled environment before facing real-world challenges. Furthermore, construction companies could adopt the methods for risk assessment, identifying potential fall risks associated with specific tasks or environments. Finally, the analysis presented in this work can serve as a tool for performance monitoring and allow timely interventions to enhance worker safety.
@@
Despite the reported benefits in terms of stability and balance, some subjects expressed that wearing the knee exoskeleton was less comfortable during kneeling compared to the stance posture. This disparity in comfort levels may be attributed to the unique biomechanical demands and joint configurations associated with kneeling. These observations underscore the importance of considering posture-specific design factors and adaptability when developing wearable exoskeletons for construction workers. There are some other limitations to consider for future research directions. First, this study only recruited young, healthy students rather than experienced construction workers, it’s important to acknowledge that experienced workers may have developed specific strategies or physical adaptations that could influence their interaction with exoskeletons. Expanding the subject pool to include experienced construction workers of varying ages will also be important to comprehensively assess the exoskeleton’s effects on balance and performance. Second, incorporating direct physiological measurements, such as electromyography (EMG), could provide deeper insights into muscle activation and neuromuscular responses. This would allow for a detailed analysis of how exoskeletons reduce physical workload on key joints. Finally, investigating fatigue effects and developing adaptive control systems that respond to real-time balance shifts would further enhance the long-term safety, safety, and efficacy of exoskeletons in construction environments.